Python 3.6

i need the code ready to copy and run with adequate indentation

Write a program that will calculate the cost of purchasing a meal. This program will include decisions and loops. Details of the program are as follows: Your menu items only include the following food with the accompanying price o Yum Yum Burger .99 o Grease Yum Fries 79 o Soda Yum 1.09 Allow the user of the program to purchase any quantity of these items in one order Allow the user of the program to purchase one or more types of these items in one order. After the order is placed, calculate total and add a 6% sales tax. Print to the screen a receipt showing the total purchase price.
Answer #1

def menu():
print("1.Yum Yum Burger = .99")
print("2.Grease Yum Fries = .79")
print("3.Soda Yum = 1.09")
print("4.Close")
i = True
j=0
k=0
l=0
total = 0
while(i):
item = int(input("Enter the no of the item "))
if(item == 1):
no = int(input("Enter no. of items u want "))
j = j+no
total = total +(no*0.99)
i = True
elif(item == 2):
noa = int(input("Enter no. of items u want "))
total = total +(noa*0.79)
k = k+noa
i = True
elif(item == 3):
now = int(input("Enter no. of items u want "))
total = total +(now*1.09)
l = l+now
i = True
elif(item == 4):
i = False
total = total+(0.06*total)

print("\n YOUR RECIEPT :")

print("1.Yum Yum Burger " + str(j) +" cost "+str(j*0.99))
print("2.Grease Yum Fries " + str(k) +" cost "+str(k*0.79))
print("3.Soda Yum " + str(l) +" cost "+str(l*1.09))
print("Total amount is " + str(total))
  

  

emint (input (Ente (item1): RESTART: F:/all/menu . ру >>import menu >>>menu.menu () 1. Yum Yum Burger = .99 2.Grease Yum Fri